assume a for-profi company has $ 8 million of long term debt with an interest rate of 6% . it has $ 3 million of preferred stock with a required divdend rate of 8% and $4 milion of common stock that is estimated to have a cost of capital of 10 % . what is its weighted average cost of capital ?
Answer #1

Total value=(8+3+4)=15 million

WACC=Respective costs*Respective weight

=(8/15*6)+(3/15*8)+(4/15*10)

=7.47%(Approx).